# Grounding or Guessing? Visual Signals for Detecting Hallucinations in Sign Language Translation

### v1

@@ 1 Introduction
Despite being a video-to-text task, sign language translation (SLT) differs fundamentally from other video understanding problems. Whereas most multimodal approaches focus on recognizing scenes or describing visual events, SLT must process a complete natural language, namely sign language, with its own vocabulary and grammar. Early approaches relied on glosses, which are explicit human annotations that map video to textual sign labels (Camgoz et al., 2018; 2020). Glosses provide strong supervision but limit scalability, since most large-scale datasets are only weakly labeled and lack gloss annotations.
@@
Hallucination detection has been studied in other multimodal settings, particularly in LVLMs and text-to-image or video generation, where strong language priors can override weaker visual signals (Chen et al., 2025; Min et al., 2025). In these domains, approaches often focus on aligning model attention with visual content or using perturbation-based analyses to identify unreliable outputs. However, SLT presents a more challenging scenario because the visual modality is not auxiliary (more than just understanding scene, detecting objects and action) but constitutes the source language itself. Consequently, hallucinations in SLT are directly equivalent to translation errors, making the detection and quantification of visual grounding critical.
@@
A crucial first step toward mitigating and reducing hallucinations is the ability to characterize and detect them, and to assign reliability scores that allow ranking candidate translations by their risk of hallucination. Text-based signals such as confidence, entropy, and perplexity capture useful linguistic uncertainty, but they miss a critical aspect in SLT: whether the output is actually supported by the signed input. We therefore introduce a complementary dimension of hallucination detection we refer to as grounding usage which quantifies the extent to which generated tokens rely on video evidence rather than language priors. By quantifying grounding usage and combining this signal with established text-only baseline measures, we can identify hallucinations arising when models guess from language model priors instead of leveraging the video stream. To this end, we define a reliability score that compares standard decoding with counterfactual runs where the video is masked or replaced. These perturbations yield two classes of evidence: (i) changes in internal states and routing (e.g., hidden-state angular shifts, cross-attention activations), and (ii) probability margins between clean and counterfactual outputs. A linear fusion of the two types produces token-level reliabilities, which are then pooled into sentence-level scores. Our experiments across datasets (PHOENIX-2014T (Camgoz et al., 2018) and CSL-Daily (Zhou et al., 2021)) and architectures (T5 (Hwang et al., 2025) and mBART (Chen et al., 2022)) demonstrate that reliability not only surpasses text-only baselines in predicting hallucinations but also generalizes robustly in cross-dataset and cross-architecture transfer. Specifically, it achieves an accuracy of 97% and maintains competitive regression performance when approximating the level of hallucination rate, with a Spearman correlation of $\rho=0.72$ (between the predicted hallucination rate and the computed hallucination rate). These findings further reveal that gloss-free models hallucinate more often because they systematically under-utilize visual information, defaulting instead to linguistic priors. Our contributions and findings can be summarized as follows:
@@
  • We introduce a novel metric based on grounding usage that quantifies how much generated tokens rely on visual input, providing a direct measure of whether SLT outputs are grounded in the source video.
@@
  • We empirically demonstrate that reliability surpasses text-only baseline metrics in detecting hallucinations and provide insights into why gloss-free models are particularly prone to them.
@@ 2.1 Sign Language Translation
Camgoz et al. (2018) proposed an end-to-end approach for generating spoken language translations directly from sign language videos and introduced the PHOENIX-2014T dataset, which provides both gloss annotations and spoken language translations for German Sign Language (DGS) videos. SLT systems and datasets are typically categorized as either gloss-based or gloss-free. Traditional SLT approaches relied on glosses as an intermediate representation. However, creating gloss annotations is time-consuming, which has led to increased interest in gloss-free SLT models. Examples of such approaches include extending Transformer architectures to SLT (Camgoz et al., 2020), leveraging sentence embeddings instead of glosses (Hamidullah et al., 2024), and integrating spatial and motion visual feature extractors with large language models (LLMs) (Hwang et al., 2025). Alongside model development, datasets for various sign languages have emerged and gained popularity, such as CSL-Daily (Zhou et al., 2021) for Chinese Sign Language (CSL) and How2Sign (Duarte et al., 2021) for American Sign Language (ASL). Using LLMs as the backbone for SLT has largely solved the fluency and scalability challenges (Wong et al., 2024; Gong et al., 2024; Inan et al., 2025), but this progress comes at a price as it increases reliance on language priors and makes the models more prone to hallucinations.
@@ 3.1 Feature-based signals: internal reliance
    $s^{\mathrm{hid}}_{t}\;=\;\pi^{-1}\arccos\!\left(\frac{\langle h^{\mathrm{vid}}_{t},\,h^{0}_{t}\rangle}{\|h^{\mathrm{vid}}_{t}\|\,\|h^{0}_{t}\|}\right),$
    where $\langle\cdot,\cdot\rangle$ denotes the inner product and $\|\cdot\|$ the Euclidean norm. A larger $s^{\mathrm{hid}}_{t}$ indicates a stronger change in orientation, and thus stronger reliance on the video input.
  • Cross-attention usage. Decoder cross-attention $A^{(\ell,h)}_{t\to i}$ at layer $\ell\in[1,L]$ and head $h\in[1,H]$ (from decoder token $t$ to encoder position $i$) is aggregated across all layers and heads. We compute the mean attention mass directed to video encoder positions and subtract the corresponding mass under a null (masked) encoder input, with per-sequence rescaling:
    $s^{\mathrm{attn}}_{t}\;=\;\mathrm{scale}\!\left(\frac{1}{LH}\sum_{\ell=1}^{L}\sum_{h=1}^{H}\sum_{i\in\mathrm{video}}A^{(\ell,h)}_{t\to i}\;-\;\frac{1}{LH}\sum_{\ell=1}^{L}\sum_{h=1}^{H}\sum_{i\in\mathrm{masked}}A^{(\ell,h)}_{t\to i}\right)$
@@ 3.2 Counterfactual signals: external evidence
$p_{\mathrm{cf}}(y_{t}\mid c_{t})\;=\;\max\big\{p_{0}(y_{t}\mid c_{t}),\;p_{\mathrm{mis}}(y_{t}\mid c_{t},x^{\prime})\big\}.$
@@ 3.3 Token-level and Sentence-level fusion of reliability signals
$r_{t}=\sigma\!\left(\mathbf{w}_{\mathrm{fb}}^{\top}\mathbf{h}_{t}\;+\;\mathbf{w}_{\mathrm{cf}}^{\top}\mathbf{g}_{t}\;+\;b\right).$
@@
We primarily rely on tail-10 pooling ($R_{\mathrm{tail}\text{-}10}$). Other options include $R_{\mathrm{mean}}$ (average over tokens), $R_{\mathrm{harm}}$ (harmonic mean), or $R_{\mathrm{min}}$ (minimum token reliability).
@@
3.4 Text only baselines and combination
@@
$\text{CHAIR}\approx\text{ISO}\big(1-R_{\ast}\big),$
@@ 4.1 Experimental Design
First, reliability features are extracted following the methodology of Section 3, with the detailed algorithm provided in Appendix D.
Next, the hallucination detection model is trained to identify hallucinated outputs, leveraging both our grounding-based reliability signals (visual features) and the baseline features described in Section 3.4, all integrated within the META (ours) framework. Accuracy is computed at a 0.5 threshold, while isotonic regression is employed to predict both $\mathrm{CHAIR}$ and its complement $(1-\mathrm{CHAIR})$. Evaluation metrics include AUROC (AUC), AUPRC (AP), and ACC@0.5 (ACC). The resulting weights are further examined for their transferability across datasets and model families.
Finally, the behavior of reliability is analyzed under both visual and linguistic degradations. Stress tests cover Gaussian noise, temporal downsampling, and random frame dropping on the visual side, as well as wrong-start prompt prefixes to assess language-model dominance during decoding. In addition to CHAIR, the analysis tracks how BLEU (Papineni et al., 2002) an $n$-gram precision metric and chrF (Popović, 2015) a character-$n$-gram F-score sensitive to morphological variation, evolve under these perturbations and how their trends correlate with reliability.
@@ 4.2 Results
Table 1 upper block shows that our grounding-based reliability consistently improves over text-only baselines on hallucination detection. Across datasets and models, we gain up to +0.20 AUC and +0.30 ACC compared to entropy and perplexity, reaching near-perfect AP ($\approx 0.99$). We present the detailed gain over the other baselines in Table 2. The META (ours) variant, which fuses grounding with text-side uncertainty signals, achieves the strongest overall detection performance, demonstrating that grounding contributes complementary information beyond text probability-based baselines. At the same time, pure Grounding (ours) is slightly better in terms of accuracy, as highlighted in the Figure 2, showing that visual evidence can sharpen decision boundaries in ways text-only signals cannot. This advantage is especially critical in sign language translation, and even more so in the gloss-free case, where no intermediate gloss supervision is available and reliable grounding in the continuous video stream is the only safeguard against hallucinations. More broadly, this underscores the need in multimodal generation to go beyond text-only uncertainty: hallucination detection benefits from explicitly checking whether outputs are supported by the input modality.
Detection
(A / P / C)
CSL
GB
CSL
GF
PHOENIX
GB
PHOENIX
GF
Confidence
A:+0.155
P:+0.015
C:+0.455
A:-0.002
P:-0.000
C:+0.102
A:+0.209
P:+0.067
C:+0.254
A:-0.008
P:-0.003
C:+0.078
Perplexity
A:+0.194
P:+0.010
C:+0.353
A:-0.001
P:-0.000
C:+0.128
A:+0.232
P:+0.062
C:+0.338
A:+0.029
P:+0.008
C:+0.224
Token entropy
A:+0.073
P:+0.005
C:+0.298
A:-0.007
P:-0.000
C:+0.105
A:+0.189
P:+0.054
C:+0.318
A:+0.047
P:+0.011
C:+0.207
@@
Regression
(P / S / I)
CSL
GB
CSL
GF
PHOENIX
GB
PHOENIX
GF
Confidence
R:-0.186
S:-0.157
I:+0.171
R:+0.003
S:-0.022
I:+0.007
R:-0.114
S:-0.092
I:+0.063
R:-0.011
S:-0.011
I:+0.013
Perplexity
R:-0.360
S:-0.344
I:+0.346
R:-0.088
S:-0.019
I:+0.003
R:-0.269
S:-0.171
I:+0.197
R:-0.223
S:-0.048
I:+0.041
Token entropy
R:-0.281
S:-0.283
I:+0.270
R:-0.012
S:+0.018
I:-0.026
R:-0.183
S:-0.127
I:+0.151
R:-0.155
S:-0.046
I:+0.025
@@
On the regression task (see Table  1 bottom block), where CHAIR is approximated as a continuous target, text-based baselines achieve slightly higher than our the correlations obtained using only our Grounding (ours) (e.g., token entropy reaches $\rho=-0.698$ on CSL GF, perplexity $\rho=-0.542$ on PHOENIX GF), reflecting their strength as direct uncertainty estimators since entropy and perplexity are tied to token-level probabilities. Our grounding signal remains competitive, with Spearman correlations up to $\rho=-0.680$ (CSL GF) and isotonic fits of 0.720.720.650.65 across datasets. Its counterfactual components exploit probability shifts between clean and perturbed video to provide an energy-based ranking that complements text baselines. The main observation is that while grounding alone does not fully match text-only uncertainty measures, it contributes complementary cues: text features capture uncertainty within the language model, while grounding verifies whether predictions are visually supported. Together, as shown by the META (ours) variant (reaching $\rho=-0.736$ on CSL GF and ISO up to $0.76$), this synergy yields the most reliable hallucination detection across datasets and transfer settings.
@@
A linear calibration trained on one dataset transfers to another one with only a small drop in regression performance, and the same holds across gloss-free $\leftrightarrow$ gloss-based models. Some loss is expected, as the calibration reflects inevitable dataset- and model-specific biases. Nevertheless, reliability remains the strongest signal across transfers (the full transfer matrix is shown in the Appendix, Fig. 8). In Fig. 3(a) we present cross-dataset transfer within the same model family (GF vs GB). It remains close, as in CSL-GF $\rightarrow$ PHOENIX-GF, where the amount of hallucinations in both in source and target might be similar and therefore the calibration signal learnable. We present the transfer between gloss-based and gloss-free models in Fig. 3(b) leads to a more pronounced degradation in performance, indicating disparities in tokenization granularity and a reduction in exposure to linguistically informative samples (as gloss-free models tend to exhibit higher hallucination rates): gloss-based systems operate mostly at the word level to maintain gloss alignment, while gloss-free systems use subunits like (Kudo & Richardson, 2018). Grounding signals align better with subunits, which clearly indicate which parts of words are guessed independently of video. Subunits level tokens often marks stems, subword boundaries which also correspond to visual units in sign language; glosses mostly represent content words, which are used as objects in CHAIR adaptation. This also suggests that transfer of the calibration to newer gloss-free architectures is more feasible, with smaller loss in terms of post-calibration performance, due to closer alignment with pretrained LLM subword tokenization. n Fig. 3(c), we also report the transfer gaps observed when transferring across both datasets and model types. The results confirm that the performance degradation is larger when reliability weights are transferred from gloss-based (GB) to gloss-free (GF) models, whereas the opposite direction (GF→GB) exhibits comparatively smaller gaps.

### v2

@@ 1 Introduction
Despite being a video-to-text task, sign language translation (SLT) differs fundamentally from other video understanding problems. Whereas most multimodal approaches focus on recognizing scenes or describing visual events, SLT must process a complete natural language, namely sign language, with its own vocabulary and grammar. Early approaches relied on glosses, which are explicit human annotations that map video to textual sign labels (Camgoz et al., 2018; 2020). Glosses provide strong supervision but limit scalability, since most large-scale SLT datasets are only weakly labeled and lack gloss annotations.
@@
Hallucination detection has been studied in other multimodal settings, particularly in LVLMs and text-to-image or video generation, where strong language priors can override weaker visual signals (Chen et al., 2025; Min et al., 2025). In these domains, approaches often focus on aligning model attention with visual content or using perturbation-based analyses to identify unreliable outputs. However, SLT presents a more challenging scenario because the visual modality is not auxiliary (more than just understanding scene, detecting objects and action) but constitutes the source language itself. Consequently, hallucinations in SLT are directly equivalent to translation errors, making the detection and quantification of visual grounding critical. In this work, we define hallucination in SLT as the generation of content tokens that are not supported by the signed video.
@@
A crucial first step toward mitigating and reducing hallucinations is the ability to characterize and detect them, and to assign reliability scores that allow ranking candidate translations by their risk of hallucination. Text-based signals such as confidence, entropy, and perplexity capture useful linguistic uncertainty, but they miss a critical aspect in SLT: whether the output is actually supported by the signed input. We therefore introduce a complementary dimension of hallucination detection we refer to as grounding usage which quantifies the extent to which generated tokens rely on video evidence rather than language priors. By quantifying grounding usage and combining this signal with established text-only baseline measures, we can identify hallucinations arising when models guess from language model priors instead of leveraging the video stream. To this end, we define a reliability score that compares standard decoding with counterfactual runs where the video is masked or replaced. These perturbations yield two classes of evidence: (i) changes in internal states and routing (e.g., hidden-state angular shifts, cross-attention activations), and (ii) probability margins between clean and counterfactual outputs. A linear fusion of the two types produces token-level reliabilities, which are then pooled into sentence-level scores. Our experiments across datasets (PHOENIX-2014T (Camgoz et al., 2018) and CSL-Daily (Zhou et al., 2021)) and architectures (T5 (Hwang et al., 2025) and mBART (Chen et al., 2022)) demonstrate that reliability not only surpasses text-only baselines in predicting hallucinations but also generalizes robustly in cross-dataset and cross-architecture transfer. Specifically, it achieves an accuracy of 97% and maintains competitive regression performance when approximating the level of hallucination rate, with a Spearman correlation of $\rho=0.72$ (between predicted and computed hallucination rates). Our findings further reveal that, compared to gloss-based models, gloss-free models hallucinate more often because they systematically under-utilize visual information, defaulting to linguistic priors. Our contributions and findings can be summarized as follows:
@@
  • We introduce a novel measure referred to as reliability based on grounding usage that quantifies how much generated tokens rely on visual input, providing a direct measure of whether SLT outputs are grounded in the source video.
@@
  • We empirically demonstrate that reliability surpasses text-only baseline measures in detecting hallucinations and provide insights into why gloss-free models are particularly prone to them.
@@ 2.1 Sign Language Translation
Camgoz et al. (2018) proposed an end-to-end approach for generating spoken language translations directly from sign language videos and introduced the PHOENIX-2014T dataset, which provides both gloss annotations and spoken language translations for German Sign Language (DGS) videos. SLT systems and datasets are typically categorized as either gloss-based or gloss-free. Traditional SLT approaches relied on glosses as an intermediate representation. However, creating gloss annotations is time-consuming, which has led to increased interest in gloss-free SLT models. Examples of such approaches include extending Transformer architectures to SLT (Camgoz et al., 2020), leveraging sentence embeddings instead of glosses (Hamidullah et al., 2024), and integrating spatial and motion visual feature extractors with large language models (LLMs) (Hwang et al., 2025). Alongside model development, datasets for various sign languages have emerged and gained popularity, such as CSL-Daily (Zhou et al., 2021) for Chinese Sign Language (CSL) and How2Sign (Duarte et al., 2021) for American Sign Language (ASL). Using LLMs as the backbone for SLT target text generation has largely solved the fluency and scalability challenges (Wong et al., 2024; Gong et al., 2024; Inan et al., 2025), but this progress comes at a price as it increases reliance on language priors and makes the models more prone to hallucinations.
@@ 3.1 Feature-based signals: internal reliance
    $s^{\mathrm{hid}}_{t}\;=\;\pi^{-1}\arccos\!\left(\frac{\langle h^{\mathrm{vid}}_{t},\,h^{0}_{t}\rangle}{\|h^{\mathrm{vid}}_{t}\|\,\|h^{0}_{t}\|}\right),$ (1)
    where $\langle\cdot,\cdot\rangle$ denotes the inner product and $\|\cdot\|$ the Euclidean norm. A larger $s^{\mathrm{hid}}_{t}$ indicates a stronger change in orientation, and thus stronger reliance on the video input. The division by $\pi$ normalizes the angular distance (in radians) to $[0,1]$ for scale consistency.
  • Cross-attention usage. Decoder cross-attention $A^{(\ell,h)}_{t\to i}$ at layer $\ell\in[1,L]$ and head $h\in[1,H]$ (from decoder token $t$ to encoder position $i$) is aggregated across all layers and heads ($L$: total decoder layers, $H$: attention heads). We compute the mean attention mass directed to video encoder positions and subtract the corresponding mass under a null (masked) encoder input, with per-sequence rescaling:
    $s^{\mathrm{attn}}_{t}\;=\;\mathrm{scale}\!\left(\frac{1}{LH}\sum_{\ell=1}^{L}\sum_{h=1}^{H}\sum_{i\in\mathrm{video}}A^{(\ell,h)}_{t\to i}\;-\;\frac{1}{LH}\sum_{\ell=1}^{L}\sum_{h=1}^{H}\sum_{i\in\mathrm{masked}}A^{(\ell,h)}_{t\to i}\right)$, (2)
    where $\mathrm{scale}(\cdot)$ applies per-sequence quantile scaling to $[0,1]$:
    A larger $s^{\mathrm{attn}}_{t}$ indicates stronger reliance on grounded (video) information.
@@ 3.2 Counterfactual signals: external evidence
$p_{\mathrm{cf}}(y_{t}\mid c_{t})\;=\;\max\big\{p_{0}(y_{t}\mid c_{t}),\;p_{\mathrm{mis}}(y_{t}\mid c_{t},x^{\prime})\big\}.$ (3)
@@ 3.3 Token-level and Sentence-level fusion of reliability signals
$r_{t}=\sigma\!\left(\mathbf{w}_{\mathrm{fb}}^{\top}\mathbf{h}_{t}\;+\;\mathbf{w}_{\mathrm{cf}}^{\top}\mathbf{g}_{t}\;+\;b\right).$ (4)
@@
where $qT$ denotes the number of tokens corresponding to the lowest $q\%$ of token reliabilities in the sequence (i.e., $\lceil qT\rceil$ tokens). We primarily rely on tail-10 pooling ($R_{\mathrm{tail}\text{-}10}$). Other options include $R_{\mathrm{mean}}$ (average over tokens), $R_{\mathrm{harm}}$ (harmonic mean), or $R_{\mathrm{min}}$ (minimum token reliability).
@@
3.4 Text only baselines and combination with grounding-based signals
@@
$\text{CHAIR}\approx\text{ISO}\big(1-R_{\ast}\big),$ (5)
@@ 4.1 Experimental Design
First, reliability features are extracted following the methodology described in Section 3, with the detailed algorithm provided in Appendix D.
Next, we train regression models for hallucination detection and for predicting CHAIR from these features, yielding reliability weights. The hallucination detection model is trained to identify hallucinated outputs, leveraging both our grounding-based reliability signals (visual features) and the baseline features described in Section 3.4, all integrated within the META (ours) framework. Accuracy is computed at a 0.5 threshold against the CHAIR-based detection label, where samples with $\mathrm{CHAIR}>0$ are labeled as 1 (hallucinated) and others as 0. Isotonic regression is employed to predict both $\mathrm{CHAIR}$ and its complement $(1-\mathrm{CHAIR})$. Evaluation metrics include AUROC (AUC), AUPRC (AP), and ACC@0.5 (ACC). The resulting weights are further examined for their transferability across datasets and model families.
Finally, reliability behavior is analyzed under both visual and linguistic degradations. Stress tests include Gaussian noise, temporal downsampling, and random frame dropping on the visual side, as well as wrong-start prompt prefixes to assess language-model dominance during decoding. In addition to CHAIR, the analysis tracks how BLEU (Papineni et al., 2002), an $n$-gram precision metric, and chrF (Popović, 2015), a character-$n$-gram F-score sensitive to morphological variation, evolve under these perturbations and how their trends correlate with reliability.
@@ 4.2 Results
Table 1 upper block shows that our grounding-based reliability consistently improves over text-only baselines on hallucination detection. Across datasets and models, we gain up to +0.20 AUC and +0.30 ACC compared to entropy and perplexity, reaching near-perfect AP ($\approx 0.99$). We present the detailed gain over the other baselines in Table 2. The META (ours) variant, which fuses grounding with text-side uncertainty signals, achieves the strongest overall detection performance, demonstrating that grounding contributes complementary information beyond text probability-based baselines. At the same time, pure Grounding (ours) is slightly better than any text-only baseline in terms of accuracy, as highlighted in the Figure 2, showing that visual evidence can sharpen decision boundaries in ways text-only signals cannot. This advantage is especially critical in sign language translation, and even more so in the gloss-free case, where no intermediate gloss supervision is available and reliable grounding in the continuous video stream is the only safeguard against hallucinations. More broadly, this underscores the need in multimodal generation to go beyond text-only uncertainty: hallucination detection benefits from explicitly checking whether outputs are supported by the input modality.
Detection (A / P / C) CSL GB CSL GF PHOENIX GB PHOENIX GF
Confidence A:+0.155 P:+0.015 C:+0.455 A:-0.002 P:-0.000 C:+0.102 A:+0.209 P:+0.067 C:+0.254 A:-0.008 P:-0.003 C:+0.078
Perplexity A:+0.194 P:+0.010 C:+0.353 A:-0.001 P:-0.000 C:+0.128 A:+0.232 P:+0.062 C:+0.338 A:+0.029 P:+0.008 C:+0.224
Token entropy A:+0.073 P:+0.005 C:+0.298 A:-0.007 P:-0.000 C:+0.105 A:+0.189 P:+0.054 C:+0.318 A:+0.047 P:+0.011 C:+0.207
@@
Regression (P / S / I) CSL GB CSL GF PHOENIX GB PHOENIX GF
Confidence R:-0.186 S:-0.157 I:+0.171 R:+0.003 S:-0.022 I:+0.007 R:-0.114 S:-0.092 I:+0.063 R:-0.011 S:-0.011 I:+0.013
Perplexity R:-0.360 S:-0.344 I:+0.346 R:-0.088 S:-0.019 I:+0.003 R:-0.269 S:-0.171 I:+0.197 R:-0.223 S:-0.048 I:+0.041
Token entropy R:-0.281 S:-0.283 I:+0.270 R:-0.012 S:+0.018 I:-0.026 R:-0.183 S:-0.127 I:+0.151 R:-0.155 S:-0.046 I:+0.025
@@
On the regression task (see Table  1 bottom block), where CHAIR is approximated as a continuous target, text-based baselines achieve slightly higher correlations compared to using only our Grounding (ours) (e.g., token entropy reaches $\rho=-0.698$ on CSL GF, perplexity $\rho=-0.542$ on PHOENIX GF), reflecting their strength as direct uncertainty estimators since entropy and perplexity are tied to token-level probabilities. Our grounding signal remains competitive, with Spearman correlations up to $\rho=-0.680$ (CSL GF) and isotonic fits of 0.720.720.650.65 across datasets. Its counterfactual components exploit probability shifts between clean and perturbed video to provide an energy-based ranking that complements text baselines. The main observation is that while grounding alone does not fully match text-only uncertainty measures, it contributes complementary cues: text features capture uncertainty within the language model, while grounding verifies whether predictions are visually supported. Together, as shown by the META (ours) variant (reaching $\rho=-0.736$ on CSL GF and ISO up to $0.76$), this synergy yields the most reliable hallucination detection across datasets and transfer settings.
@@
A linear calibration trained on one dataset transfers to another one with only a small drop in regression performance, and the same holds across gloss-free $\leftrightarrow$ gloss-based models. Some loss is expected, as the calibration reflects inevitable dataset- and model-specific biases. Nevertheless, reliability remains the strongest signal across transfers (the full transfer matrix is shown in the Appendix, Fig. 8). In Fig. 3(a) we present cross-dataset transfer within the same model family (GF vs GB). It remains close, as in CSL-GF $\rightarrow$ PHOENIX-GF, where the amount of hallucinations in both in source and target might be similar and therefore the calibration signal learnable. As presented in Fig. 3(b), the transfer between gloss-based and gloss-free models leads to a more pronounced degradation in performance, indicating disparities in tokenization granularity and a reduction in exposure to linguistically informative samples (as gloss-free models tend to exhibit higher hallucination rates): gloss-based systems operate mostly at the word level to maintain gloss alignment, while gloss-free systems use subword units (Kudo & Richardson, 2018). Grounding signals align better with subunits, which clearly indicate which parts of words are guessed independently of video. Subunits level tokens often mark stems, subword boundaries which also correspond to visual units in sign language; glosses mostly represent content words, which are used as objects in CHAIR adaptation. This also suggests that transfer of the calibration to newer gloss-free architectures is more feasible, with smaller loss in terms of post-calibration performance, due to closer alignment with pretrained LLM subword tokenization. In Fig. 3(c), we also report the transfer gaps observed when transferring across both datasets and model types. The results confirm that the performance degradation is larger when reliability weights are transferred from gloss-based (GB) to gloss-free (GF) models, whereas the opposite direction (GF→GB) exhibits comparatively smaller gaps.